# A Neural Network Enhanced Born Approximation for Inverse Scattering

### v2

@@ 3.1 Discretization of the Born approximation
We assume a standard source-receiver setup. The incident and scattered fields are sampled uniformly in $\mathbb{S}^{1}$, namely,
@@
where $N_{S}$ and $N_{R}$ denote the number of sources and receivers, respectively. For simplicity, we consider the case in which $N_{S}=N_{R}$, following [10]. Let $\hat{\textbf{x}}_{i}=\langle\cos\theta_{i},\sin\theta_{i}\rangle$ and $\textbf{d}_{j}=\langle\cos\varphi_{j},\sin\varphi_{j}\rangle$. By varying both the incident and scattered field directions, it follows that we have available the far field matrix $U_{\infty}\in\mathbb{C}^{N_{R}\times N_{S}}$ possibly corrupted by noise. Using quadrature on (11), we obtain the components for the discrete Born approximation for a single incident field
$(U_{\infty})_{i,j}\approx\frac{k^{2}}{\sqrt{8\pi k}}\exp\left(\frac{\mathrm{i}\mkern 1.0mu\pi}{4}\right)h^{2}\sum_{p,q=1}^{n}\exp(\mathrm{i}\mkern 1.0muk(\textbf{d}_{j}-\hat{\textbf{x}}_{i})\cdot\textbf{y}_{p,q})M_{p,q},$ (13)
where $h=2/N_{g}$. The above approximation in tensor form can be written as a product between the Born $4$-tensor $\mathscr{B}_{\rm disc}\in\mathbb{C}^{N_{R}\times N_{S}\times N_{g}\times N_{g}}$ and the scatterer matrix $M$
@@
where $\mathscr{B}_{\rm disc}$ encodes the (2+2)-discretization in both angular dimensions and both spatial dimensions, and $\tau_{\mu}$ is the unknown error for the Born approximation.
To utilize linear algebraic methods, we can use standard tensor unfolding methods to collapse $\mathscr{B}_{\rm disc}$ along the spatial and angular dimensions to produce the aforementioned stacked Born operator $\mathcal{B}\in\mathbb{C}^{N_{R}\times N_{S}\times N_{g}^{2}}$. We thus arrive at the following formulation:
@@
where $\gamma>0$ is a regularization parameter. Experimentally, we observe that an optimal choice of $\gamma$ will generally be in the range $[0.1,1]$ for the inverse problems in our study.
@@ 3.2 Data generation
As discussed in Section 3.1, we utilize synthetic scattering data generated using a standard finite element approach to approximating (2). For our experiments, we fix the wave number at $k=16$ and set the number of sources and receivers to be $N_{S}=N_{R}=100$. Higher wave number problems may also be learned, provided $N_{S}$ and $N_{R}$ are increased. In the case of training data, we do not add extra measurement noise to the computed far field matrix $U_{\infty}$. For a discussion of generalization to the case with noise, see Section 5.2. We also assume that no absorption occurs, i.e., $\Im(\eta(\textbf{x}))=0$. For a discussion of generalization to the case with weak, random absorption, see Section 5.3. The spatial discretization of the domain $\Omega=[-1,1]^{2}$ uses $N_{g}=100$.
@@
For training, we want to use simple scatterers. Here we create the scatterer field $\eta(\textbf{x})$ as the union of $N_{c}$ piecewise-constant circles overlaying a homogeneous background of air, where we uniformly sample $N_{c}\in\{1,2,3\}$. The $i$th circle is assigned a constant value $\eta_{i}$ sampled from the uniform distribution $\mathcal{U}(1.1,1.8)$, with the background air set to $\eta=1$ (we shall see that $\eta\approx 1.8$ is outside the weak scattering regime). The radius and position of each circle are sampled uniformly from $\mathcal{U}(0.1,0.3)$ and $\mathcal{U}(-0.7,0.7)$, respectively. In selecting the training data, for simplicity of mesh generation, we enforce that the circles cannot overlap and must be fully contained in the search domain $[-1,1]^{2}$. We generate 20,000 samples for training and validation (under an 80-20% split) with an additional 4,000 samples for testing purposes. Visualizations of typical training samples can be seen in Figure 2.
The validity of the Born approximation is contingent upon weak scattering of the incident field. Moreover, the convergence of the Neumann series (and hence the accuracy of the Born approximation) requires $\eta$ to be sufficiently close to unity. In the case of the synthetic dataset, the Born approximation fails when $\eta_{i}\gg 1$ for any circle, thus preventing an accurate inversion of (14). In particular, a direct inversion results in severe underestimates of the true contrast regardless of the chosen regularization parameter $\gamma$.
When weak scattering breaks down, the Born approximation also gives rise to artifacts when there is multiple scattering. Examples of poor reconstructions using the Born approximation are shown in Section 5. We seek to remediate these failures in the strong scattering case through corrective approaches using neural networks.
@@ 4.1 Correction strategies
We propose two methods for correcting the Born approximation: a pre-correction and a post-correction. Let $\vec{\mu}_{\rm exact}$ denote the exact nodal values for the (vectorized) contrast. There is an error $\vec{\tau}_{\mu}$ associated with both the convergence of the Neumann series and its truncation (the Born approximation). We may write
@@
where $\mathcal{B}$ is the stacked Born operator defined in (15), and $\vec{u}_{\infty}$ is the exact far field pattern. The factor $\vec{\tau}_{\mu}$ can be thought of as the discrepancy in the Born-obtained far field compared to the true (simulated) far field. A CNN is trained to predict $\vec{\tau}_{\mu}$, and this approach pre-corrects the far field data into a form suitable for the Born approximation to predict an accurate contrast. In particular
@@
provided that $\gamma$ is chosen correctly and $\vec{u}_{\infty}^{\tau_{\mu}}$ is learned appropriately. This is motivated by the approach in [15], where a neural network is employed to compensate for modeling errors introduced by approximate forward models.
@@ 4.2 Training
For the two correction strategies, we can associate the following CNN-Born (CNNB) and Born-CNN (BCNN) models, respectively. The former performs the Born approximation on the corrected input far field while the latter performs the Born approximation on the labels of the training set. For a training set of size $N_{\rm train}$ we have the corresponding training regimes:
  • CNNB: We consider data pairs $\{(\vec{u}_{\infty}^{(i)},\vec{u}^{\tau_{m}(i)}_{\infty})\}_{i=1}^{N_{\rm train}}$ and seek to minimize the loss function
@@
  • BCNN: We consider data pairs $\{(\vec{u}_{\infty}^{(i)},\varepsilon_{\mu}^{(i))}\}_{i=1}^{N_{\rm train}}$ and seek to minimize the loss function
@@
    where $\rm BCNN(\>\cdot\>;\beta^{\prime})$ is another CNN weights and biases collected in the vector $\beta^{\prime}$.
To provide a baseline for comparison, we train a simple black-box CNN to directly map the far field data to the solution of the inverse scattering problem motivated by [10]. In other words, we consider data-pairs $\{(\vec{u}_{\infty}^{(i)},\vec{\mu}^{(i)})\}_{i=1}^{N_{\rm train}}$ and minimize the loss function
@@
where $\rm CNN(\>\cdot\>,\beta^{\prime\prime})$ is a third CNN with weights and biases $\beta^{\prime\prime}$.
To aid comparison we use the same general CNN structure in all three cases, but optimize the network hyper-parameters to each case. Furthermore, we compute the basic regularized Born inverse (16) with regularization parameter $\gamma=1$ and $\gamma=0.1$ for a direct comparison of CNN results to the Born approximation.
[…]
Figure 3: The neural network architecture for the three models compared in this paper. (a) The CNN model directly maps the far field to the squared refractive index. (b) The regularized Born inverse is computed to obtain a rough estimate of the squared refractive index $\eta(\textbf{x})$, which is then corrected by the BCNN predicted $\varepsilon$. (c) The CNNB model pre-corrects the far field and then applies a regularized Born inverse to obtain the squared refractive index. Note that the necessary reshaping of the far field are omitted in this diagram, as well as the map $\mu\mapsto\mu+1=\eta$.
In all three approaches, the model takes as input a representation of the far field matrix, specifically, where the real and imaginary parts are separated and stored in two distinct channels, which enables the model to process both components simultaneously. The model output is then reshaped into a two-dimensional map corresponding to the discrepancy term. This end-to-end mapping from far field data to an image-like discrepancy allows the model to leverage spatial correlations in the data that are challenging to capture through purely analytic approaches. The structure of the three methods is visualized in Figure 3.
@@
The architecture of each CNN was determined through a randomized search process, where 150 network configurations were generated, and the best-performing model was selected based on the lowest validation mean squared error (MSE). The randomization process involved varying hyper-parameters, including the number of convolution layers (ranging from 1 to 4) and the number of fully connected (FC) layers (ranging from 1 to 3). The number of channels per each convolution layer was randomly chosen between 16 and 512, while the number of units per FC layer was selected between 64 and 1024. Max pooling, with a fixed kernel size of 2, was applied after each convolutional layer.
For activation functions, we randomly selected from a set including Rectified Linear Unit (ReLU), LeakyReLU with a negative slope of 0.1, Gaussian Error Linear Unit (GELU), Sin, and Sigmoid. The last layer always used a linear activation function. All models were trained using the AdamW optimizer with an initial learning rate of 0.0005, adjusted via a learning rate scheduler that applied a 5% decay every 100 epochs. The models were trained with an early stopping patience of 50 epochs. Table 1 provides a summary of the tuned hyper-parameters.
@@ 5.1 In-distribution performance
Following the discussion in Section 3.1, we generated 4,000 in-distribution scatterers to evaluate model performance. Table 2 summarizes the average $\ell^{2}$ and $\ell^{1}$ errors of the test scatterer reconstructions, and Figure 4 shows the actual profile of one example of a test scatterer together with the reconstructions.
To make clear that the CNN based models outperform the inverse Born approximation by itself, Figure 5 visualizes the error distribution for the test samples. The results show that the three models–CNN, CNNB, and BCNN–significantly outperform the Born approximation in terms of accuracy, reducing the relative $\ell^{2}$ error by almost 50%.
The distribution of relative errors reveals that the CNN based models consistently produce low-error reconstructions, with a tightly clustered distribution around the mean, compared to the direct Born approximation. This suggests that the learning-based models generalize well within the in-distribution setting. In contrast, the Born approximation exhibits a broader and more skewed error distribution, highlighting its poor accuracy in reconstructing scatterers outside of the weak-scattering approximation.
Table 2: Average relative $\ell^{2}$ and $\ell^{1}$ error in the reconstruction of the contrast $\mu$ expressed as a percentage on the test dataset of $4,000$ scatterers. The rows labeled Born show the results of using the regularized inverse Born approximation. The CNNB model performs optimally in both $\ell^{2}$ and $\ell^{1}$.
$\ell^{2}$ error (%) $\ell^{1}$ error (%)
@@
Figure 4: An example of the reconstructions of three circular scatterers determined by each model. Here we show density plots of the exact and reconstructed values of $\eta$. The ground truth is shown in the upper left panel. All CNN based methods are much more accurate than the inverse Born approximation since, in this example, the scattering is strong. CNNB (bottom left) shows the least speckle in the background.
Figure 5: The distribution of relative $\ell^{2}$ (left) and $\ell^{1}$ (right) norm errors for the test dataset. The average of each distribution is denoted by the corresponding colored circle. The CNN based methods all improve over the simple Born approximation.
@@ 5.2 Noise robustness
Noise, arising from practical measurement apparatus errors or environmental factors, often corrupts scattering data, posing significant challenges to accurately reconstructing scatterers. We trained the CNNs using noise free data (apart from numerical error), so it is important to determine how the various CNN models cope with out-of-distribution data corresponding to far field patterns corrupted by noise. In this context, following [10], we consider a far field matrix affected by noise, with entries given by
@@
where $\delta$ controls the strength of the noise and $\mathcal{E}_{i,j}$ is the (complex-valued) noise for incoming wave direction $\varphi_{i}$ and measurement direction $\theta_{j}$. To model the noise, we assume that $\mathcal{E}_{i,j}$ is sampled from a univariate complex standard normal distribution ${\cal CN}(0,1)$ [1, Def 2.1].

### v3

@@ 3.1 Discretization of the Born approximation
We assume a standard source-receiver setup. The incident and scattered fields are uniformly in $\mathbb{S}^{1}$, using angles
@@
where $N_{S}$ and $N_{R}$ denote the number of sources and receivers, respectively. For simplicity, we consider the case in which $N_{S}=N_{R}$, following [15]. Let $\hat{\textbf{x}}_{i}=\langle\cos\theta_{i},\sin\theta_{i}\rangle$ and $\textbf{d}_{j}=\langle\cos\varphi_{j},\sin\varphi_{j}\rangle$. By varying both the incident and scattered field directions, it follows that we have available the far field matrix $U_{\infty}\in\mathbb{C}^{N_{R}\times N_{S}}$ possibly corrupted by measurement noise or forward modeling errors.
Using quadrature on (11), we obtain the components for the discrete Born approximation for a single incident field
$(U_{\infty})_{i,j}\approx\frac{k^{2}}{\sqrt{8\pi k}}\exp\left(\frac{\mathrm{i}\mkern 1.0mu\pi}{4}\right)h^{2}\sum_{p,q=0}^{N_{g}}\exp(\mathrm{i}\mkern 1.0muk(\textbf{d}_{j}-\hat{\textbf{x}}_{i})\cdot\textbf{y}_{p,q})M_{p,q},$ (13)
where $h=2/N_{g}$. The above approximation in tensor form can be written as a product between the Born 444-tensor $\mathscr{B}_{\rm disc}\in\mathbb{C}^{N_{R}\times N_{S}\times N_{g}\times N_{g}}$ and the scatterer matrix $M$
@@
where $\mathscr{B}_{\rm disc}$ encodes the (2+2)-discretization in both angular dimensions and both spatial dimensions, and $\tau_{\mu}$ is the unknown error due to the Born approximation.
To utilize linear algebraic methods, we can use standard tensor unfolding methods to collapse $\mathscr{B}_{\rm disc}$ along the spatial and angular dimensions to produce the aforementioned stacked Born operator $\mathcal{B}\in\mathbb{C}^{N_{R}N_{S}\times N_{g}^{2}}$. We thus arrive at the following formulation:
@@
where $\gamma>0$ is a regularization parameter. Experimentally, we observe that an optimal choice of $\gamma$ will generally be in the range [0.1,1][0.1,1][ 0.1 , 1 ] for the inverse problems in our study.
@@ 3.2 Data generation
As discussed at the end of Section 3.1, we utilize synthetic scattering data generated using a standard finite element approach to approximating (2). For our experiments, we fix the wave number at $k=16$ as in [15] and set the number of sources and receivers to be $N_{S}=N_{R}=100$. Higher wave number problems may also be learned, provided $N_{S}$ and $N_{R}$ are increased. In the case of training data, we do not add extra measurement noise to the computed far field matrix $U_{\infty}$. For a discussion of generalization to the case with noise, see Section 5.2. For our training data, we also assume that no absorption occurs, i.e., $\Im(\eta(\textbf{x}))=0$. For a discussion of generalization to the case with weak absorption, see Section 5.3. The spatial discretization of the domain $\Omega=[-1,1]^{2}$ uses $N_{g}=100$.
For training, we want to use simple scatterers. Here we create the scatterer field $\eta(\textbf{x})$ as the union of $N_{c}$ piecewise-constant circles overlaying a homogeneous background of air, where we uniformly sample $N_{c}\in\{1,2,3\}$. The $i$th circle is assigned a constant value $\eta_{i}$ sampled from the uniform distribution $\mathcal{U}(1.1,1.8)$, with the background air set to $\eta=1$ (we shall see that $\eta\approx 1.8$ is outside the weak scattering regime). The radius and position of each circle are sampled uniformly from $\mathcal{U}(0.1,0.3)$ and $\mathcal{U}(-0.7,0.7)$, respectively. In selecting the training data, for simplicity of mesh generation, we enforce that the circles cannot overlap and must be fully contained in the search domain $[-1,1]^{2}$. We generate 20,000 samples for training and validation (under an 80-20% split) with an additional 4,000 samples for testing purposes. This choice is motivated by [15] where 16,384 samples are used for training. Visualizations of typical training samples can be seen in Fig. 2.
The validity of the Born approximation is contingent upon weak scattering of the incident field. Moreover, the convergence of the Neumann series (and hence the accuracy of the Born approximation) requires $\eta$ to be sufficiently close to unity. In the case of the synthetic dataset, the Born approximation fails when $\eta_{i}\gg 1$ for any circle, thus preventing an accurate inversion of (14). In particular, a direct inversion results in severe underestimates of the true contrast regardless of the chosen regularization parameter $\gamma$. When weak scattering breaks down, the Born approximation also gives rise to artifacts when there is multiple scattering. Examples of poor reconstructions using the Born approximation are shown in Section 5. We seek to remediate these failures in the strong scattering case through corrective approaches using NNs.
@@ 4.1 Correction strategies
We suggest two methods for correcting the Born approximation: a pre-correction and a post-correction. Let $\vec{\mu}_{\rm exact}$ denote the exact nodal values for the (vectorized) contrast. There is an error $\vec{\tau}_{\mu}$ associated with both the convergence of the Neumann series and its truncation (the Born approximation). We may write
@@
where $\mathcal{B}$ is the stacked Born operator defined in (15), and $\vec{u}_{\infty}$ is the exact far field pattern. The factor $\vec{\tau}_{\mu}$ can be thought of as the discrepancy in the Born-obtained far field compared to the true far field. A CNN is trained to predict $\vec{\tau}_{\mu}$, and this approach pre-corrects the far field data into a form suitable for the Born approximation to predict an accurate contrast. In particular
@@
provided that $\gamma$ is chosen correctly and $\vec{u}_{\infty}^{\tau_{\mu}}$ is learned appropriately. This is motivated by the approach in [23], where a NN is employed to compensate for modeling errors introduced by approximate forward models.
@@ 4.2 Training
[…]
Figure 3: The neural network architecture for the three models compared in this paper. (a) The CNN, NIO, and BCR models directly map the far field to the squared refractive index. Note that the BCR model only uses $\Re(\vec{u}_{\infty})$ as input. (b) The regularized Born inverse is computed to obtain a rough estimate of the squared refractive index $\eta(\textbf{x})$, which is then corrected by the BCNN predicted $\varepsilon$. (c) The CNNB model pre-corrects the far field and then applies a regularized Born inverse to obtain the squared refractive index. Note that the necessary reshaping of the far field and contrast are omitted in this diagram, as well as the map from the contrast to the refractive index: $\mu\mapsto\mu+1=\eta$.
For the two correction strategies, we can associate the following CNN-Born (CNNB) and Born-CNN (BCNN) models, respectively (see Fig. 3). The former performs the Born approximation on the corrected input far field while the latter performs the Born approximation on the labels of the training set. For a training set of size $N_{\rm train}$ we have the corresponding training regimes:
  • CNNB: We consider data pairs $\{(\vec{u}_{\infty}^{(i)},\vec{u}^{\tau_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu}}(i)}_{\infty})\}_{i=1}^{N_{\rm train}}$ and seek to learn $\vec{\tau}_{\mu}$ by minimizing the loss function
@@
  • BCNN: We consider data pairs $\{(\vec{u}_{\infty}^{(i)},\varepsilon_{\mu}^{(i))}\}_{i=1}^{N_{\rm train}}$ and seek to learn $\vec{\varepsilon}_{\mu}$ by minimizing the loss function
@@
    where $\rm BCNN(\>\cdot\>;\beta^{\prime})$ is another CNN with weights and biases collected in the vector $\beta^{\prime}$.
To provide a baseline for comparison, we train a simple black-box CNN to directly map the far field data to the solution of the inverse scattering problem motivated by [15]. In other words, we consider data-pairs $\{(\vec{u}_{\infty}^{(i)},\vec{\mu}^{(i)})\}_{i=1}^{N_{\rm train}}$ and seek to directly learn $\vec{\mu}$ by minimizing the loss function
@@
where $\rm CNN(\>\cdot\>,\beta^{\prime\prime})$ is a third CNN with weights and biases $\beta^{\prime\prime}$. To aid comparison we use the same general CNN structure in all three cases, but optimize the network hyper-parameters to each case. Furthermore, we compute the basic regularized Born inverse (16) with regularization parameter $\gamma=1$ and $\gamma=0.1$ for a direct comparison of CNN results to the Born approximation.
@@
Besides the straight-foward CNN, we also compare to results computed using the implementation of NIO [31] available from Github [^1] https://github.com/mroberto166/nio. Although there is a Helmholtz-based example in [31], it is not for scattering on an infinite domain (rather it is for Dirichlet and Neumann data pairs on a bounded domain). Nonetheless, the software can be easily modified to handle far field data from incident plane waves as in our case. We trained NIO using our data as discussed previously for CNN with an $\ell_{2}$ loss function. Using scripts from the GitHub repository, we optimized NIO for our problem.
We also compare to the BCR-Net of [14, 15], trained using our data with $\ell^{2}$ loss and utilizing the training scripts and network code generously provided to us by its authors. Training on our data is necessary to make a comparison with our networks meaningful. In the CNN and NIO approaches, the model takes as input a representation of the far field matrix, where the real and imaginary parts are separated and stored in two distinct channels, which enables the model to process both components simultaneously. Meanwhile, the BCR-Net model takes solely the real part of the far field matrix as input. The model output is then reshaped into a two-dimensional map corresponding to the discrepancy term. This end-to-end mapping allows the model to leverage spatial correlations in the data that are challenging to capture through purely analytic approaches. The high-level structure of the CNN based methods is visualized in Fig.3. For the structure of BCR-Net see [15] and for NIO see [31].
The architecture of each CNN was determined through a randomized search process, where 150 network configurations were generated, and the best-performing model was selected based on the lowest validation mean squared error (MSE). The randomization process involved varying hyper-parameters, including the number of convolution layers (ranging from 1 to 4) and the number of fully connected (FC) layers (ranging from 1 to 3). The number of channels per each convolution layer was randomly chosen between 16 and 512, while the number of units per FC layer was selected between 64 and 1024. Max pooling, with a fixed kernel size of 2, was applied after each convolutional layer. For activation functions, we randomly selected from a set including Rectified Linear Unit (ReLU), LeakyReLU with a negative slope of 0.1, Gaussian Error Linear Unit (GELU), Sine, and Sigmoid. The last layer always used a linear activation function. All CNN based models were trained using the AdamW optimizer (discussed in [29]) with an initial learning rate of 0.0005, adjusted via a learning rate scheduler that applied a 5% decay every 100 epochs. The models were trained with an early stopping patience of 50 epochs. Table 1 provides a summary of the tuned hyper-parameters.
@@ 5.1 In-distribution performance
Following the methods discussed in Section 3.1, we generated 4,000 in-distribution scatterers to evaluate model performance. Table 2 summarizes the average $\ell^{2}$ and $\ell^{1}$ errors of the test scatterer reconstructions. Table 2 shows that while NIO and BCR-Net slightly outperform the CNN based models in both norms, all NN based models perform similarly well in $\ell^{2}$ (training used this norm for the loss function). The results also show that all NN based models significantly outperform the Born approximation in terms of accuracy, reducing the relative $\ell^{2}$ error by almost 50%. This is likely because the Born approximation performs poorly for large contrast scatterers which are included in the testing data.
To investigate further how the NN based models compare to the inverse Born approximation by itself, Fig. 4 visualizes the error distribution for the test samples. The results show that all NN based models significantly outperform the basic inverse Born approximation in terms of accuracy, reducing the relative $\ell^{2}$ error by almost 50%. This is likely because the Born approximation performs poorly for large contrast scatterers which are included in the testing data. Fig 5 shows the actual profile of one test scatterer together with the reconstructions. Qualitatively, we observe that NIO and BCR-Net sometimes have a tendency to forgot smaller, weaker scatterers. Going forward, we will omit the error and reconstruction results related to the inverse Born approximation for $\gamma=0.1$ due to its consistently sub-par performance in relation to when $\gamma=1$.
Table 2: Average relative $\ell^{2}$ and $\ell^{1}$ error (%) in the reconstruction of the contrast $\mu$ on the test dataset of 4,0004,0004 , 000 scatterers. The best reconstruction for each metric is in boldface.
Method $\ell^{2}$ (%) $\ell^{1}$ (%)
@@
Figure 4: The distribution of relative $\ell^{2}$ (left) and $\ell^{1}$ (right) norm errors for the test dataset. The average of each distribution is denoted by the corresponding colored circle. The NN based methods all improve over the simple Born approximation.
Figure 5: A sample reconstruction of three in distribution circular scatterers using each model. All NN based methods are much more accurate than the Born approximation since, in this example, the scattering is strong. While NIO (III) and BCR (IV) produce sharper edges of the scatterer, in this example they completely fail to detect the smaller weak scatterer.
@@ 5.2 Noise robustness
Figure 6: Visualization of a scatterer and its noisy far-field patterns for multiple noise levels in correspondence with (22). The left column shows the refractive index of the scatterer. The top and bottom rows display the real and imaginary parts of the far-field data for $\delta=0.1,0.25,0.5,1.0$, respectively. The reconstruction of this scatterer is shown in Fig. 7.
Noise, arising from practical measurement apparatus errors or environmental factors, often corrupts scattering data, posing significant challenges to accurately reconstructing scatterers. We trained each model using noise free data (apart from numerical error), so it is important to determine how the various models cope with out-of-distribution data corresponding to noisy far field data. In this context, following [15], we consider a far field matrix $U^{\delta}_{\infty}$ affected by noise, with entries given by
@@
where $\delta$ controls the strength of the noise and $\mathcal{E}_{i,j}$ is the (complex-valued) noise for incoming wave direction $\varphi_{i}$ and measurement direction $\theta_{j}$. To model the noise, we assume that $\mathcal{E}_{i,j}$ is sampled from a uni-variate complex standard normal distribution ${\cal CN}(0,1)$ [1, Def 2.1].
Table 3: Average relative $\ell^{2}$ and $\ell^{1}$ errors (%) in the reconstruction of the contrast $\mu$ on the test dataset of 4000 scatterers with noise added to the far field data having parameter $\delta=0.1,0.25,0.5,1.0$ defined in (22). The best reconstruction for each $\delta$ is in boldface.
$\delta=0.1$ $\delta=0.25$
Method $\ell^{2}$ (%) $\ell^{1}$ (%) Method $\ell^{2}$ (%) $\ell^{1}$ (%)
Born ($\gamma=1$) 67.0378 111.8617 Born ($\gamma=1$) 67.0378 111.8617
NIO 36.8788 45.1435 NIO 63.1058 160.1682
BCR-Net 33.8625 33.8459 BCR-Net 34.8135 34.8141
CNN 36.1031 78.3238 CNN 41.9604 90.9867
CNNB 35.8799 65.9740 CNNB 39.2874 79.4563
BCNN 38.4137 102.7461 BCNN 48.4455 143.9246
$\delta=0.5$ $\delta=1.0$
Method $\ell^{2}$ (%) $\ell^{1}$ (%) Method $\ell^{2}$ (%) $\ell^{1}$ (%)
Born ($\gamma=1$) 67.0378 111.8617 Born ($\gamma=1$) 67.0378 111.8617
NIO 103.4922 358.3474 NIO 148.9648 535.9080
BCR-Net 37.8709 38.2524 BCR-Net 47.3537 51.6489
CNN 62.6698 135.3118 CNN 134.2359 354.2700
CNNB 49.2495 116.5801 CNNB 73.5768 183.3327
BCNN 89.3354 281.7972 BCNN 380.3745 1173.8297
We analyze the zero-shot performance of the models on the testing dataset described previously under varying levels of added random noise using $\delta=0.1$, 0.250.250.25, 0.50.50.5, 111. We also compare their reconstructions to the Born model ($\gamma=1$), which is particularly robust to noise due to the regularized inversion, yet under-approximates the contrast for strong scatterers. Table 3 provides statistics on the behavior of the models.
At all levels of noise, Table 3 shows that BCR-Net is the most stable to noise. All NN based schemes perform roughly equivalently for lower noise levels at $\delta=0.1,0.25$ with CNNB being the most stable of the Born based techniques. The standard Born approximation with $\gamma=1$ is strongly regularized to noise, but cannot handle large contrasts resulting in poor average performance. We conjecture that the Born inversion incorporated after the CNN in CNNB smooths high frequency error effectively controlling noise amplification and making it the more stable of the Born based techniques. In contrast, BCNN first provides a smooth reconstruction, and then applies the CNN using that reconstruction and the far field data to derive a corrected model. This could result in increased sensitivity to noise since CNN is used after regularization. Fig. 7 shows the results for one reconstruction at various noise levels. In this case the ground truth scatterer is shown in Fig. 6. At high noise levels, this example shows that both NIO and BCR-Net may hallucinate as might be expected [17]. BCNN and CNN do not perform as well as CNNB at the highest noise level.
Figure 7: The refractive index estimated by each model corresponding to the scatterer in Fig. 6 with noisy data. In this example, the high noise regime, NIO and BCR tend to hallucinate additional scatterers while BCNN and CNN are marred with artifacts. Meanwhile, CNNB demonstrates relative robustness even at $\delta=1$ without becoming highly corrupted by background artifacts.
5.3 Absorption
Figure 8: Visualization of an absorbing scatterer for multiple absorption levels in correspondence with 23. The left column shows the real part of the scatterer with $\eta_{0}=1.8$. The right column displays the imaginary part of the scatterer for $\alpha=0,0.1,1.0,10.0$, respectively. The reconstruction of the scatterer is shown in Fig. 9
Our models are trained on data from absorption free media ($\Im(\eta)=0$), as would be appropriate for applications involving non-dissipative media. In many practical applications, the media exhibits varying degrees of absorption [19], characterized by $\Im(\eta)>0$. Of course, the inverse Born approximation handles absorbing media seamlessly (provided the scattering is weak), but the NN based schemes depend on the training data.
Since we did not train the NN based models using far field data from absorbing scatterers, it is important to determine if the NN based models can predict the real part of $\eta$ even when the scattering media are slightly absorbing.
Table 4: Average relative $\ell^{2}$ and $\ell^{1}$ errors (%) in the reconstruction of the contrast $\mu$ on the test dataset of 100 absorbing scatterers with $\alpha=0,0.1,1,10$ following (23). The best reconstruction for each $\alpha$ is in boldface.
$\alpha=0$ $\alpha=0.1$
Method $\ell^{2}$ (%) $\ell^{1}$ (%) Method $\ell^{2}$ (%) $\ell^{1}$ (%)
Born ($\gamma=1$) 88.9285 144.9036 Born ($\gamma=1$) 89.4273 133.2401
NIO 32.5428 36.7441 NIO 33.3716 38.7768
BCR-Net 41.2125 45.9028 BCR-Net 42.7416 49.2050
CNN 37.7394 78.1480 CNN 39.9897 77.3409
CNNB 37.4499 65.0130 CNNB 38.2855 68.8976
BCNN 41.6169 107.8156 BCNN 43.1420 103.9886
$\alpha=1$ $\alpha=10$
Method $\ell^{2}$ (%) $\ell^{1}$ (%) Method $\ell^{2}$ (%) $\ell^{1}$ (%)
Born ($\gamma=1$) 96.2615 119.4021 Born ($\gamma=1$) 95.1858 147.9940
NIO 67.1786 83.2703 NIO 94.7267 144.64945
BCR-Net 61.9917 85.0740 BCR-Net 72.26803 117.4409
CNN 71.2832 109.7257 CNN 78.2193 137.1136
CNNB 70.4011 107.2159 CNNB 77.0885 136.1209
BCNN 70.9577 139.9843 BCNN 78.3788 188.4496
To investigate this scenario, we consider the case in which $\eta$ inside a circular scatterer is perturbed by an unknown, small absorption term. Specifically, we use the following complex-valued $\eta$ to generate far field test data:
$\eta^{\alpha}\mathrel{\mathop{:}}=\eta_{0}+\mathrm{i}\mkern 1.0mu\alpha,$ (23)